OpenStudy (anonymous):

$25,000 is invested at 4% compounded annually for 7 years, find the current balance in the bank account based on the given conditions

OpenStudy (anonymous):

its for liberal arts

zepdrix (zepdrix):

The formula for finding compound interest is as follows,\[\large A=P\left(1+\frac{r}{n}\right)^{nt}\] They gave us this information, ~our `principle amount` \(P\) is $25,000. ~the `rate of compounding` \(r\) is 4%, or as a decimal, 0.04. ~the `time period` \(t\) is 7 years. ~the `number of compoundings per period` \(n\) is 1 (since it's compounded anually). So we just plug these values in and solve for \(A\)! :)

zepdrix (zepdrix):

\[\large A=25000\left(1+\frac{0.04}{1}\right)^{1\times7}\]

zepdrix (zepdrix):

You'll want to use your calculator to simplify that down. Lemme know if it's too complicated :D